# Thai Large Language Models

### v1

@@ Abstract
This paper introduces Typhoon 2, a series of text and multimodal large language models optimized for the Thai language. The series includes models for text, vision, and audio. Typhoon2-Text builds on state-of-the-art open models, such as Llama 3 and Qwen2, and we perform continual pre-training on a mixture of English and Thai data. We employ various post-training techniques to enhance Thai language performance while preserving the base models’ original capabilities. We release text models across a range of sizes, from 1 to 70 billion parameters, available in both base and instruction-tuned variants. Typhoon2-Vision improves Thai document understanding while retaining general visual capabilities, such as image captioning. Typhoon2-Audio introduces an end-to-end speech-to-speech model architecture capable of processing audio, speech, and text inputs and generating both text and speech outputs simultaneously.
@@
| Model | Base Model | Link |
@@ 1 Introduction
In South East Asia (SEA), the efforts to enhance foundation language models for SEA languages include SeaLLM (Zhang et al., 2024), SEA-LION (Singapore, 2024), and Sailor (Dou et al., 2024). When it comes to the Thai language, there are model families such as OpenThaiGPT (Yuenyong et al., 2024), Typhoon (Pipatanakul et al., 2023), WangChan (Polpanumas et al., 2023), and Pathumma (NECTEC, 2024).
@@ 2.1.1 Base Corpus Preparation
Scaling the Data: We initiate the process by increasing the number of Common Crawl (CommonCrawl) packages compared to the previous iteration, where we process a total of 40 packs of the CommonCrawl data.
Text Extraction: Although CommonCrawl provides a pre-extracted WET subset, several studies suggest that WET files exhibit lower quality compared to WARC files extracted using external tools such as Trafilatura (Li et al., 2024a; Penedo et al., 2023). In our study, we begin with a total of 40 packs of Thai CommonCrawl data with a cut-off date of September 2023 and perform HTML extraction from scratch using Trafilatura222https://trafilatura.readthedocs.io/en/latest/. This process results in a significantly larger dataset, comprising approximately 3 TB of Thai text, or approximately 200 billion Llama3 tokens.
Strict Deduplication: To ensure data quality and minimize redundancy, we implement a fuzzy deduplication pipeline using MinHash and LSH algorithms.
Heuristic Filtering: To filter out pages containing excessive search engine optimization (SEO) content, very short documents, and low-quality texts, we evaluate each line using signals such as the ratio of numbers to text and the punctuation density. At the document level, document filtering is done based on other metrics, including newline ratios and overall document length. After deduplication and heuristic filtering, we have approximately 44B tokens of Thai text.
@@ 2.2 Gathering Diverse and High-Quality Thai Documents
As a result, we augment our dataset with additional 12B high-quality Thai tokens through this process.
@@ 2.2.4 High-Educational Content
To gather high educational content in Thai, we use a similar approach as in collected culture-related text in Thai and fine-web edu (Penedo et al., 2024). Specifically, using an LLM to annotate 50K samples and fine-tune the BGE-M3 classification head. We filter only with a predicted value of 3 or higher due to the low level of educational content.
@@ 2.5 Evaluation
We evaluate the models using the ThaiExam and M3Exam datasets. While this evaluation method has been widely used for pre-trained language models, the scores obtained using these datasets are highly above the average level of a typical Thai person333https://www.niets.or.th/th/content/view/11821. This can be attributed to contamination and saturation due to overfitting (Fourrier et al., 2024). Nevertheless, we utilize this signal to measure whether the model has acquired knowledge of the Thai language and context, as well as a development signal for improvement.
@@ 2.6 Results and Findings
The evaluation results are shown in Table 2. We found that following insights from our experiments and evaluations.
Each filtering subset has its own performance gain: We observe a performance gain in the “Science" score on “High-Quality Text Selection," while we achieve a “Social" score on “Culture Related Text in Thai" and a “Thai" score on the “General subset" during the first-stage data mixture setup on M3Exam.
@@ 3.1.1 Data
English Instruction-Dataset: We combine several public datasets based on multiple iterations of our Typhoon development. In this version, we incorporate SystemChat444https://huggingface.co/datasets/abacusai/SystemChat-1.1, Capybara555https://huggingface.co/datasets/LDJnr/Capybara, OpenChat (Wang et al., 2024a) and a subset of the Tulu 3 (Lambert et al., 2024) dataset to retain the base model’s English language comprehension. These datasets are selected primarily to align with human feedback patterns and various use cases in leveraging LLMs. For example, SystemChat supports system role-following features; Capybara facilitates multi-turn conversations, and OpenChat and Tulu 3 provide a rich diversity of instructions.
@@
TyphoonIF Dataset: A new dataset is constructed based on AutoIF (Dong et al., 2024), available in both Thai and English. Seed constraints for the English and Thai versions are manually crafted to represent typical prompts that humans might use when interacting with LLMs in specific tasks. Additionally, random queries are taken from airesearch/WangchanThaiInstruct (Vistec, 2024) and Suraponn/thai_instruction_sft666https://huggingface.co/datasets/Suraponn/thai_instruction_sft, along with randomly sampled queries from the English dataset as previously described in this subsection. Rejected samples, identified through an instruction evaluation function, are filtered and added to the final dataset. The final instruction set comprises 150,000 question/instruction and response pairs. These pairs are randomly designated as either system or user turns to enhance generalization. To further improve cross-lingual transfer capabilities, instruction turns are randomly translated between Thai and English encouraging the models to align both Thai and English spaces to be similar to each other.
@@ 3.1.2 Experimental Setup
For IFEval and code-switching (CS) evaluation, the metric is accuracy, defined as:
@@ 3.3.1 Data
Anti-haystack777https://huggingface.co/datasets/wenbopan/anti-haystack: This dataset is designed for enhancing the ability of LLMs to locate short and precise facts within long, noisy documents, emulating a “needle in a haystack" challenge.
@@ 3.3.3 Results and Findings
Our CPT approach was restricted to a context length of 8,192 tokens, potentially limiting the model’s ability to generalize to longer contexts, despite adjustments to the RoPE scaling (Su et al., 2023).
@@
Based on our experiments, a good dataset composition consists of 40% long-context data (with a 20:80 ratio of Thai to English) and 60% short-context data. This approach yields the best overall performance across both long- and short-context tasks.
@@ 3.4.1 Data
Glaive-v2888https://huggingface.co/datasets/glaiveai/glaive-function-calling-v2: The popular Glaive-function-calling-v2 dataset is incorporated into the training data mix, complementing APIGen and ToolACE datasets to provide a comprehensive foundation for training.
@@ 3.5 Distillation
Distillation is a widely recognized method to transfer knowledge from a stronger model to a smaller model, as seen from the era of BERT to DistilBERT, and so on. In the case of LLMs, there have also been successful examples such as Minitron (Sreenivas et al., 2024) and Llama 3.2999https://ai.meta.com/blog/llama-3-2-connect-2024-vision-edge-mobile-devices/. Following this approach, we apply a distillation technique to enhance the performance of the smaller Typhoon models.
@@ 3.5.1 Top-k Logits Distillation
We employ top-k logits distillation, drawing inspiration from Arcee-AI’s methodology101010https://blog.arcee.ai/introducing-arcee-supernova-medius-a-14b-model-that-rivals-a-70b-2. The objective of this approach is to transfer knowledge from a larger teacher model to a smaller student model. Our method involves constructing logit data from a larger teacher model, where only the top-k predictions for each vocabulary token are retained. The top-k logits are obtained from early versions of the Llama-based Typhoon 2 models, specifically the 8B and 70B variants. This method assumes that the teacher and student models share identical vocabularies.
@@ 3.5.3 Results and Findings
Impact of Distillation on Performance: Based on the results, distillation yields a performance improvement in most of the aspects for the small model.
Does a larger model’s logits improve performance? In our preliminary experiments, we distil the logits from both Typhoon2-8B and Typhoon2-70B models. Based on this setup, we found that using logits from the 8B model and the 70B model result in a similar performance level.
@@ 3.6 Model Merging
Model merging, a method to combine the weights of two models, has recently been shown to improve performance in LLMs (Akiba et al., 2024) . We previously performed model merging for our Typhoon 1.5X series111111https://blog.opentyphoon.ai/typhoon-1-5x-our-experiment-designed-for-application-use-cases-7b85d9e9845c, resulting in significant improvements for Thai and English instruction-following tasks. Notably, our larger-size models, such as those with 70B parameters, demonstrated remarkable performance gains. Therefore, we apply model merging techniques to our 70B models in this iteration.
@@ 4.3 Thai OCR Enhancement
In our efforts to improve the intricacies and detailed analysis within the domains of the Thai Economic Report-Finance Research and the Thai Book, we deploy Optical Character Recognition (OCR) methodologies. By incorporating the advanced capabilities of Agentics, we further augment and refine the ground-truth labelling processes that were initially derived from the comprehensive dataset in its entirety.
@@ 4.3.1 Data Sources: Thai Book and Thai Financial
As shown in Table 12, our goal is on the intricate process of data filtration, which serves to enhance and elevate the caliber of data employed in our comprehensive analysis. Initially, the dataset includes a substantial volume of images and text dialogues obtained from the Thai Economic Report and the Thai Book. Through our multi-step filtering methodologies, the dataset has been reduced to preserve only the entries that are deemed most pertinent and exhibit the highest quality, deemed essential for subsequent analytical processing.
@@ 4.3.2 Agentic Framework for Thai VQA Data Curation
In dominian data-centric AI (DCAI) emerges as a paradigm aimed at improving the veracity label of the data sets. This is achieved by applying an agentic framework that systematically revisits the perspective through which data are analysed and interpreted, thereby refining the truthfulness inherent within the data. The method is tailored to tackle highly intricate tasks, beginning with initial labels and adjusting them to accurately represent the data.
@@
We adhered to the outlined automated design framework specific to agentic systems as documented in Hu et al. (2024) for structuring each individual agent. Thereafter, our method was influenced by the Buffer of Thoughts (BoT) approach detailed in Yang et al. (2024c), incorporating the technique of routing within the meta-prompt as a central feature. Throughout the development stages of each generational cycle, our methodology closely mirrors the strategy observed in Close-Quarters Battle (CQB), focusing on pinpointing areas that require refinement and ensuring appropriate modifications are executed. Additionally, this system facilitates the transfer of knowledge across different iterations, making it somewhat analogous to the ToT framework while maintaining a maximum depth level of 3. Our method distinctly diverges as it includes a mechanism for assigning a self-evaluation score aimed at reducing self-bias, a principle discussed in Xu et al. (2024).
@@ 2. Processing via Agentic Refine Steps
The ultimate content chosen from Agentic, as indicated by Section 4.3, has been used to produce data sets consisting of question-context-answer dialogues. This specific Agentic data set was generated using the application of LlamaIndex. Subsequent refinement was achieved by integrating the Synthesiser-Refine technique into the process. Furthermore, our meta-prompt-CoT strategy was used to systematically create questions, relying exclusively on the capabilities of the Typhoon-1.5X model.
The methodologies discussed previously were similarly implemented on an open-source external dataset. In these cases, the foundational approach was preserved; however, specific modifications were introduced. These alterations included adjusting the meta-prompt identification number and, in certain situations, opting to employ a singular agent rather than the complete multi-agent configuration.
Use Cases of Data Filtered
Utilizing the Typhoon 1.5X model, we refined the content pattern by excluding refusals from all images. This process employed the sentence transformers paraphrase-multilingual-MiniLM-L12-v2 model with a similarity threshold of 0.2337.
In the subsequent phase of our research, we have meticulously compiled a comprehensive dataset specifically tailored for fine-tuning instruction-based tasks conducted in the Thai language. This dataset is instrumental for the automation of issue detection and labeling. These functions are critical elements in the methodology aimed at transitioning towards the establishment of a reliable and trustworthy model.
@@ 4.4 Experimental Setup
This evaluation examined the performance enhancements of tasks such as ChartQA, MTVQ (TH), OCR (TH), OCRBench, and TextVQA against baseline models. We concentrated on advancements in Thai-specific OCR and visual question answering. Key metrics evaluated accuracy, efficiency, and model robustness across these tasks. The findings indicated substantial improvements in our model over the base models, especially in challenging scenarios like ChartQA and MTVQ. This evaluation underscores the model’s effectiveness in tackling both general and domain-specific issues.
@@ 4.5 Results and Findings
In our instruction tuning experiments, we consider two base models, Llama 3.2 and Qwen2, and we employ the same data set to fine-tune both models. Their performance results are detailed in Table 13. Following these evaluations, Typhoon2-qwen2vl-7b-vision-instruct is chosen for our release due to its superior performance and computational efficiency. We balance the representation of Thai and English tasks by selecting 4 metrics per language, to ensure fairness and thoroughness in our study. The selection of Typhoon2-Qwen2-VL is influenced by these factors:
Efficiency: Despite having fewer parameters, the model meets the performance standards effectively.
Superior Performance: It excels in key areas such as ChartQA, OCR (TH), MTVQ (TH), and M3Exam Images (TH).
ChartQA Emphasis: ChartQA took precedence due to inadequate existing solutions (e.g., Docling, EasyOCR, PyTesseract), which focus primarily on TextVQA and do not adequately address the unique challenges of ChartQA.
However, the similar-performing Typhoon2-llama-3.2-11B-Instruct (Exp) is not chosen. This was due to the priority of task-specific strengths and the real-world applicability of Typhoon2-qwen2vl-7b-vision-instruct being prioritised.
| Benchmark | Llama-3.2 | Qwen2-VL | Pathumma | Typhoon2-llama-3.2 | Typhoon2-qwen2vl |
| 11B-Instruct | 7B-Instruct | Vision-1.0.0-8B | 11B-Instruct (Exp) | 7B-vision-instruct | |
| OCRBench | 72.84 | 72.31 | 32.74 | 81.20 | 64.38 |
| Liu et al. (2024c) | 51.10 | 57.90 | 25.87 | 71.70 | 49.60 |
| MMBench (Dev) | - | - | - | - | - |
| Liu et al. (2024b) | 76.54 | 84.10 | 19.51 | 83.66 | 83.66 |
| ChartQA | 13.41 | 47.45 | 64.20 | 74.12 | 75.71 |
| Masry et al. (2022) | x | 45.00 | 57.83 | 67.36 | 72.56 |
| TextVQA | 32.82 | 91.40 | 32.54 | 89.44 | 91.45 |
| Singh et al. (2019) | x | 88.70 | 28.84 | 85.74 | 88.97 |
| OCR (TH) | 64.41 | 56.47 | 6.38 | 79.51 | 64.24 |
| Sapsathien & Jaroenkantasima (2024) | 35.58 | 55.34 | 2.88 | 58.65 | 63.11 |
| M3Exam Images-(TH) | - | - | - | - | - |
| Zhang et al. (2023c) | 25.46 | 32.17 | 29.01 | 27.93 | 33.67 |
| GQA (TH) | 31.33 | 34.55 | 10.20 | 44.51 | 50.25 |
| Hudson & Manning (2019) | - | - | - | - | - |
| MTVQ (TH) | 11.21 | 23.39 | 7.63 | 15.20 | 30.59 |
| Tang et al. (2024b) | 4.31 | 13.79 | 1.72 | 7.56 | 21.55 |
| Average (Upper) | 37.67 | 54.26 | 25.61 | 64.16 | 62.77 |
| Average (Lower) | x | 53.85 | 23.67 | 58.75 | 59.02 |

### v2

@@ Abstract
This paper introduces Typhoon 2, a series of text and multimodal large language models optimized for the Thai language. The series includes models for text, vision, and audio. Typhoon2-Text builds on state-of-the-art open models, such as Llama 3 and Qwen2, and we perform continual pre-training on a mixture of English and Thai data. We employ post-training techniques to enhance Thai language performance while preserving the base models’ original capabilities. We release text models across a range of sizes, from 1 to 70 billion parameters, available in both base and instruction-tuned variants. To guardrail text generation, we release Typhoon2-Safety, a classifier enhanced for Thai cultures and language. Typhoon2-Vision improves Thai document understanding while retaining general visual capabilities, such as image captioning. Typhoon2-Audio introduces an end-to-end speech-to-speech model architecture capable of processing audio, speech, and text inputs and generating both text and speech outputs.
@@
| Model | Base Model | Link to HuggingFace |
@@ 1 Introduction
In South East Asia (SEA), the efforts to enhance foundation language models for SEA languages include SeaLLM (Zhang et al., 2024), SEA-LION (Singapore, 2024), and Sailor (Dou et al., 2024). When it comes to the Thai language, there are model families such as WangChan (Polpanumas et al., 2023), Typhoon (Pipatanakul et al., 2023), OpenThaiGPT (Yuenyong et al., 2024), and Pathumma (NECTEC, 2024).
@@ 2.1.1 Base Corpus Preparation
Step1 - Scaling the Data: We initiate the process by increasing the number of Common Crawl (CommonCrawl) packages compared to the previous iteration, where we process a total of 40 packs of the CommonCrawl data.
Step2 - Text Extraction: Although CommonCrawl provides a pre-extracted WET subset, several studies suggest that WET files exhibit lower quality compared to WARC files extracted using external tools such as Trafilatura (Li et al., 2024a; Penedo et al., 2023). In our study, we begin with a total of 40 packs of Thai CommonCrawl data with a cut-off date of September 2023 and perform HTML extraction from scratch using Trafilatura222https://trafilatura.readthedocs.io/en/latest/. This process results in a significantly larger dataset, comprising approximately 3 TB of Thai text, or approximately 200 billion Llama3 tokens.
Step3 - Strict Deduplication: To ensure data quality and minimize redundancy, we implement a fuzzy deduplication pipeline using MinHash and locality-sensitive hashing (LSH) algorithms.333https://github.com/ChenghaoMou/text-dedup
Step4 - Heuristic Filtering: To filter out pages containing excessive search engine optimization (SEO) content, very short documents, and low-quality texts, we evaluate each line using signals such as the ratio of numbers to text and the punctuation density. At the document level, document filtering is done based on other metrics, including newline ratios and overall document length. After deduplication and heuristic filtering, we have approximately 44B tokens of Thai text.
@@ 2.2 Gathering Diverse and High-Quality Thai Documents
As a result, we augment our dataset with an additional 12B high-quality Thai tokens through this process.
@@ 2.2.4 High-Educational Content
To gather high educational content in Thai, we use a similar approach as in collected culture-related text in Thai and fine-web edu (Penedo et al., 2024). Specifically, an LLM is used to annotate 50K samples and we fine-tune the BGE-M3 classification head. We filter only with a predicted value of 3 or higher due to the low level of educational content.
@@ 2.5 Evaluation
We evaluate the models using the ThaiExam and M3Exam datasets. While this evaluation method has been widely used for pre-trained language models, the scores obtained using these datasets are highly above the average level of a typical Thai person444https://www.niets.or.th/th/content/view/11821. This can be attributed to contamination and saturation due to overfitting (Fourrier et al., 2024). Nevertheless, we utilize this signal to measure whether the model has acquired knowledge of the Thai language and context, as well as a development signal for improvement.
@@ 2.6 Results and Findings
The evaluation results are shown in Table 2. We found the following insights from our experiments and evaluations.
Each filtering subset has its own performance gain: We observe a performance gain in the “Science" score on “High-Quality Text Selection," while we achieve a gain in the “Social" score on “Culture Related Text in Thai" and a gain in the “Thai" score on the “General subset" during the first-stage data mixture setup on M3Exam.
@@ 3.1.1 Data
English Instruction-Dataset: We combine several public datasets based on multiple iterations of our Typhoon development. In this version, we incorporate SystemChat555https://huggingface.co/datasets/abacusai/SystemChat-1.1, Capybara666https://huggingface.co/datasets/LDJnr/Capybara, OpenChat (Wang et al., 2024a) and a subset of the Tulu 3 (Lambert et al., 2024) dataset as examples to retain the base model’s English language comprehension. These datasets are selected primarily to align with human feedback patterns and various use cases in leveraging LLMs. For example, SystemChat supports system role-following features; Capybara facilitates multi-turn conversations, and OpenChat and Tulu 3 provide a rich diversity of instructions.
@@
TyphoonIF Dataset: A new dataset is constructed based on AutoIF (Dong et al., 2024), available in both Thai and English. Seed constraints for the English and Thai versions are manually crafted to represent typical prompts that humans might use when interacting with LLMs in specific tasks. Additionally, random queries are taken from airesearch/WangchanThaiInstruct (Vistec, 2024) and Suraponn/thai_instruction_sft777https://huggingface.co/datasets/Suraponn/thai_instruction_sft, along with randomly sampled queries from the English dataset as previously described in this subsection. Rejected samples, identified through an instruction evaluation function, are filtered and added to the final dataset. The final instruction set comprises 150,000 question/instruction and response pairs. These pairs are randomly designated as either system or user turns to enhance generalization. To further improve cross-lingual transfer capabilities, instruction turns are randomly translated between Thai and English encouraging the models to align both Thai and English spaces to be similar to each other.
@@ 3.1.2 Experimental Setup
For Code-switching (CS) evaluation, the metric is accuracy, defined as:
@@ 3.3.1 Data
Anti-haystack888https://huggingface.co/datasets/wenbopan/anti-haystack: This dataset is designed for enhancing the ability of LLMs to locate short and precise facts within long, noisy documents, emulating a “needle in a haystack" challenge.
@@ 3.3.3 Results and Findings
Our CPT approach is restricted to a context length of 8,192 tokens, potentially limiting the model’s ability to generalize to longer contexts, despite adjustments to the RoPE scaling (Su et al., 2023).
@@
Based on our experiments, a good dataset composition consists of 15% long-context data (with a 20:80 ratio of Thai to English) and 85% short-context data. This approach yields the best overall performance across both long- and short-context tasks.
@@ 3.4.1 Data
Glaive-v2999https://huggingface.co/datasets/glaiveai/glaive-function-calling-v2: The popular Glaive-function-calling-v2 dataset is incorporated into the training data mix, complementing APIGen and ToolACE datasets to provide a comprehensive foundation for training.
@@ 3.5 Distillation
Distillation is a widely recognized method to transfer knowledge from a stronger model to a smaller model, as seen from the era of BERT to DistilBERT(Sanh et al., 2020), and so on. In the case of LLMs, there have also been successful examples such as Minitron (Sreenivas et al., 2024) and Llama 3.2101010https://ai.meta.com/blog/llama-3-2-connect-2024-vision-edge-mobile-devices/. Following this approach, we apply a distillation technique to enhance the performance of the smaller Typhoon models.
@@ 3.5.1 Top-k Logits Distillation
We employ top-k logits distillation, drawing inspiration from Arcee-AI’s methodology111111https://blog.arcee.ai/introducing-arcee-supernova-medius-a-14b-model-that-rivals-a-70b-2. The objective of this approach is to transfer knowledge from a larger teacher model to a smaller student model. Our method involves constructing logit data from a larger teacher model, where only the top-k predictions for each vocabulary token are retained. The top-k logits are obtained from early versions of the Llama-based Typhoon 2 models, specifically the 8B and 70B variants. This method assumes that the teacher and student models share identical vocabularies.
@@ 3.5.3 Results and Findings
Impact of Distillation on Performance: Based on the results, distillation yields a performance improvement in most of the aspects of the small model.
Does a larger model’s logits improve performance? In our preliminary experiments, we distil the logits from both Typhoon2-8B and Typhoon2-70B models. Based on this setup, we observe a similar results on the 1B and 3B models in our settings.
@@ 3.6 Model Merging
Model merging, a method to combine the weights of two models, has recently been shown to improve performance in LLMs (Akiba et al., 2024) . We previously performed model merging for our Typhoon 1.5X series121212https://blog.opentyphoon.ai/typhoon-1-5x-our-experiment-designed-for-application-use-cases-7b85d9e9845c, resulting in significant improvements for Thai and English instruction-following tasks. Notably, our larger-size models, such as those with 70B parameters, demonstrated remarkable performance gains. Therefore, we apply model merging techniques to our 70B models in this iteration.
@@ 4.3 Thai OCR Enhancement
In our efforts to enhance the capabilities of document understanding for Thai Vision, with a focus on finance and general Thai books, we present our systematic Agentics methods to improve and enable tasks such as ChatQA, Document VQA, Data Visualisation QA, and Image Captioning within documents. This pipeline is designed to follow a structured and systematic approach:
1.
Data Preparation: Initially, data from the Thai Economic Report-Finance Research and the Thai Book are collected and organised in Section 4.3.1.
2.
Content Extraction: Open-weight models are used to derive structured and unstructured content from this data set.
3.
Agentic Refine Ground Truth and TextVQA Generation: The extracted data undergo agentic refinement to improve the quality of the label and annotation, while Agentic systems simultaneously generate TextVQA data, including tasks like Caption, Q&A, and Conversations, as explained in detail in Section 4.3.
@@ 4.3.1 Data Sources: Thai Book and Thai Financial
As shown in Table 12, our goal is on the intricate process of data filtration, which serves to enhance and elevate the caliber of data employed in our comprehensive analysis. Initially, the data set includes a substantial volume of images and text dialogues obtained from the Thai Economic Report and the Thai Book. Our multi-step filtering methodologies (to be described in Section 4.3.2) are designed to filter out data and retain only the entries that are deemed most pertinent and exhibit the highest quality, deemed essential for subsequent analytical processing.
@@ 4.3.2 Agentic Framework for Thai VQA Data Curation
In dominian data-centric AI (DCAI), it emerges as a paradigm aimed at improving the veracity label of the data sets. This is achieved by applying an agentic framework that systematically revisits the perspective through which data are analysed and interpreted, thereby refining the truthfulness inherent within the data. The method is tailored to tackle highly intricate tasks, beginning with initial labels and adjusting them to accurately represent the data.
@@
We adhere to the outlined automated design framework specific to agentic systems as documented in Hu et al. (2024) for structuring each individual agent. Thereafter, our method is influenced by the Buffer of Thoughts (BoT) approach (Yang et al., 2024c), incorporating the technique of routing within the meta-prompt as a central feature. Throughout the development stages of each generational cycle, our methodology closely mirrors the strategy observed in Close-Quarters Battle (CQB), focusing on pinpointing areas that require refinement and ensuring appropriate modifications are executed. Additionally, this system facilitates the transfer of knowledge across different iterations, making it similar to the ToT framework while maintaining a maximum depth level of 3. Our method distinctly diverges as it includes a mechanism for assigning a self-evaluation score aimed at reducing self-bias, a principle discussed in Xu et al. (2024).
@@ 2. Processing via Agentic Refine Steps
The ultimate content chosen from Agentic Refine Ground Truth, as indicated by Section 4.3, has been used to produce data sets consisting of question-context-answer dialogues. This specific Agentic data set is generated using LlamaIndex. Subsequent refinement is achieved by integrating the Synthesiser-Refine technique into the process. Furthermore, our meta-prompt-CoT strategy is used to systematically create questions, relying exclusively on the capabilities of the Typhoon-1.5X model.
The methodologies discussed previously are implemented using an open-source external dataset. In these cases, the foundational approach is preserved; however, specific modifications are introduced. These alterations include adjusting the meta-prompt identification number and, in certain situations, opting to employ a singular agent rather than the complete multi-agent configuration.
Data Filtering
Utilizing the Typhoon 1.5X model, we refine the content pattern by excluding question-answer pairs that have no answers from data. This process employs the sentence transformers paraphrase-multilingual-MiniLM-L12-v2 model with a similarity threshold of 0.2337.
In the subsequent phase of this experiment, we compile a comprehensive dataset specifically tailored for fine-tuning instruction-based tasks conducted in the Thai language. This dataset is instrumental in the automation of issue detection and labeling. These functions are critical elements in the methodology aimed at transitioning towards the establishment of a reliable and trustworthy model.
@@ 4.4 Experimental Setup
This evaluation evaluates models’ performance in tasks such as ChartQA, MTVQ (TH), OCR (TH), OCRBench and TextVQA against other baselines, including Llama-3.2-Vision, Qwen2-VL, and Pathumma-Vision. We focus on Thai-specific OCR and visual question-answering tasks. We use standard evaluation metrics for these tasks which are Accuracy and ROUGE-L – the measure of the longest n-gram overlap between the generated text and its reference.
@@ 4.5 Results and Findings
In our instruction tuning experiments, we consider two base models, Llama 3.2 and Qwen2, and employ the same data set to fine-tune both models.
We evaluate the performance in both Thai and English tasks using four evaluation datasets per language. The results in Table 13 show the following findings:
Efficiency: Typhoon2-Qwen2-VL, despite have fewer parameters, can match the performance of Typhoon2-Llama-3.2.
Superior Performance: Typhoon2-Qwen2-VL excels in key areas such as ChartQA, OCR (TH), MTVQ (TH), and M3Exam Images (TH) compared to other competitive models.
ChartQA Emphasis: ChartQA takes precedence due to inadequate existing solutions (e.g., Docling, EasyOCR, PyTesseract), which focus primarily on TextVQA and do not adequately address the unique challenges of ChartQA.
Based on these results, Typhoon2-Qwen2-VL is chosen for this release due to its superior performance and fewer parameter counts.
| Benchmark | Metric | Llama-3.2 | Qwen2-VL | Pathumma | Typhoon2-llama-3.2 | Typhoon2-qwen2vl |
| 11B-Instruct | 7B-Instruct | Vision-1.0.0-8B | 11B-Instruct (Exp) | 7B-vision-instruct | ||
| OCRBench | ROUGE-L | 72.84 | 72.31 | 32.74 | 81.20 | 64.38 |
| Liu et al. (2024c) | Accuracy | 51.10 | 57.90 | 25.87 | 71.70 | 49.60 |
| MMBench (Dev) | ROUGE-L | - | - | - | - | - |
| Liu et al. (2024b) | Accuracy | 76.54 | 84.10 | 19.51 | 83.66 | 83.66 |
| ChartQA | ROUGE-L | 13.41 | 47.45 | 64.20 | 74.12 | 75.71 |
| Masry et al. (2022) | Accuracy | x | 45.00 | 57.83 | 67.36 | 72.56 |
| TextVQA | ROUGE-L | 32.82 | 91.40 | 32.54 | 89.44 | 91.45 |
| Singh et al. (2019) | Accuracy | x | 88.70 | 28.84 | 85.74 | 88.97 |
| OCR (TH) | ROUGE-L | 64.41 | 56.47 | 6.38 | 79.51 | 64.24 |
| Sapsathien & Jaroenkantasima (2024) | Accuracy | 35.58 | 55.34 | 2.88 | 58.65 | 63.11 |
| M3Exam Images-(TH) | ROUGE-L | - | - | - | - | - |
| Zhang et al. (2023c) | Accuracy | 25.46 | 32.17 | 29.01 | 27.93 | 33.67 |
| GQA (TH) | ROUGE-L | 31.33 | 34.55 | 10.20 | 44.51 | 50.25 |
| Hudson & Manning (2019) | Accuracy | - | - | - | - | - |
| MTVQ (TH) | ROUGE-L | 11.21 | 23.39 | 7.63 | 15.20 | 30.59 |
| Tang et al. (2024b) | Accuracy | 4.31 | 13.79 | 1.72 | 7.56 | 21.55 |
| Average (ROUGE-L) | 37.67 | 54.26 | 25.61 | 64.16 | 62.77 | |
| Average (Accuracy) | x | 53.85 | 23.67 | 58.75 | 59.02 |